I don't know what a theoretical probability is. The question is: Megan rolls a number cube whose faces are labeled 7 through 12. Find the probability of P(even) and write it as a fraction. Please help me.
Answer by ewatrrr(24785)

 
Hi,
Megan rolls a number cube whose faces are labeled 7 through 12.
7, 8, 9, 10, 11, 12 6 faces total ( 3 of which are even numbers)
P(even) = 3/6 = 1/2